OpenStudy (anonymous):

In her first five basketball games, Sarah scored 2, 5, 10, 15, and 18 points. What is the mean absolute deviation for this set of scores? I got 5.2 points. Am I right?

OpenStudy (anonymous):

@iGreen @perl

OpenStudy (love_ranaa):

Three steps: 1. Find the mean of all values 2. Find the distance of each value from that mean (subtract the mean from each value, ignore minus signs) 3. Then find the mean of those distances That too much and thats why I hate math :)

OpenStudy (anonymous):

2 + 5 + 10 + 15 + 18 = 50 50 / 5 = 10 2 - 10 = -8 5 - 10 = -5 10 - 10 = 0 15 - 10 = 5 18 - 10 = 8 8 + 5 + 0 + 5 + 8 = 26 0 doesn't count really o-o and you could just take away the -'s So the MAD Would be 26. LOL This MAD has way to many steps.......

OpenStudy (anonymous):

@Love_Ranaa

OpenStudy (love_ranaa):

But this is my first time doing this and I got 5.2 too ! YASSS...I mean, sure I think its 5.2 ;)

OpenStudy (anonymous):

26 / 5 = 5.2 so 5.2 is exactly the MAD

OpenStudy (love_ranaa):

And IKR. Wayyy too many steps

OpenStudy (anonymous):

LOL @Love_Ranaa Thanks boo ;]

OpenStudy (love_ranaa):

Yeah So you're right :)

OpenStudy (love_ranaa):

No problem :) And my answer shows up late since this website is slow

OpenStudy (anonymous):

Same here ;-; lagsalot.com

OpenStudy (love_ranaa):

IKR but if you need help again, just tag me :)